The sum of digits of a 2 digit number is 10.The digits of the number interchange their places, when 36 is substracted from the number.Find the two digit number
Answer by sachi(548)

The sum of digits of a 2 digit number is 10.
let the ones digit=x then tenth digit=10-x
the no is 10[10-x]+x
when the digits interchange their position the no will be 10x+[10-x]
The digits of the number interchange their places, when 36 is subtracted from the number.
so 10x+[10-x]=10[10-x]+x-36
or 10x+10-x=100-10x+x-36
or 18x=54
or x=3
so the two digit number=10[10-x]+x=10[10-3]+3=73